OpenStudy (anonymous):

Solve the first order IVP. dy/dx +P(x)y=x^-6; y(1)=-1 P(x)={5/x, 1<_x<_e} P(x)={-1/x, x>e}

OpenStudy (dumbcow):

For 1st condition: 1<=x<e e^int(P(x)) = x^5 x^5(dy/dx) + 5x^4y = x^-1 (x^5y)' = x^-1 integrate both sides x^5y = ln(x) +C y = ln(x)/x^5 + C y(1) = -1 -1 = ln(1)/1 + C -1 = C y = (ln(x)/x^5) -1

OpenStudy (dumbcow):

not sure if you need to work out 2nd condition since initial value has already been satisfied

OpenStudy (anonymous):

How did you get ***=ln(x)+c? Shouldn't it be -1/5 x^-5 +c?

OpenStudy (dumbcow):

oh i multiplied the x^-6 by x^5 resulting in x^-1

OpenStudy (dumbcow):

im pretty sure you multiply the x^5 on both sides to keep equation balanced

OpenStudy (dumbcow):

oh i multiplied the x^-6 by x^5 resulting in x^-1

OpenStudy (anonymous):

Dang it, I always forget that. Is that the correct form for the equation for both of them, because I remember my professor saying that you have to do something differently with a piecewise equation. I can't remember what though.

OpenStudy (dumbcow):

yeah thats what im not sure on either...sorry

OpenStudy (anonymous):

Thanks anyway. Maybe I'kk post it again later...

OpenStudy (anonymous):

Anyone else know?

OpenStudy (anonymous):

IIRC you take the iv for the second piece by by plugging in the end of domain of first piece which is start of second (x=e?) The second piece is a different equation. Your result will be piecewise as well...ln etc, 1<_x<_e and whatever, x>e.

OpenStudy (anonymous):

I'm really not sure I understand what you're saying. Can you show me with the problem?

OpenStudy (anonymous):

You got y = (ln(x)/x^5) -1 for first piece, 1<_x<_e. Plug in e (end of domain for first piece, start of domain for second piece) to get ln(e)/e^5 -1 = e^-5-1 and this is your "y1" for the second leg where (P(x) = -1/x).